# SDP bounds on quantum codes

### v1

@@ 8.2. Stabilizer codes
Consider the special case of a stabilizer code with stabilizer group $S$. Then we can recover the results in Propositions 16 and 17 in an alternative fashion for the matrices
@@
that is, by replacing $Q=\{0^{n}\}$ with $Q=\mathcal{S}$ where $\mathcal{S}$ contains the unsigned Pauli strings from $S$ mapped to $4^{n}$.
Then as in the classical case discussed in Sections 6.3 and 6.4, these matrices are both positive semidefinite and invariant under $\operatorname{Aut}_{0}(4,n)$ and can thus be expressed in a basis of the Terwilliger algebra $\mathcal{A}_{4,n}$, recovering the same constraints on $x^{t,p}_{i,j}$ as in the general case. This follows from the fact that by the definition of $\Gamma^{v}$ in Eq. (94), it holds that $\Gamma^{v}=\Gamma^{0}$ for any $v\in\mathcal{S}$ leading to $\tilde{\Gamma}=\Gamma^{0}$, and that for a stabilizer code $|S|=|\mathcal{S}|=2^{n}/K$.
@@
which recovers the general case [see Eq. (109)]. In analogy to Proposition (17), the expansion in the Terwilliger algebra reads
@@
(133) $\tilde{\Gamma}^{c}(\mathcal{S})$ $=\frac{|\mathcal{S}|}{4^{n}-|\mathcal{S}|}\sum_{(i,j,t,p)\in\mathcal{I}(q,n)}(x^{0,0}_{i+j-t-p}-x^{t,p}_{i,j})M^{t,p}_{i,j}\,,$
@@ 8.3. Symmetry-reduced constraints
Before starting with the proof, it is useful to recall from Eq. (124) that
@@
(i) From $\Gamma_{00}=1$, $(M_{0,0}^{0,0})_{00}=1$ and zero otherwise, and $\gamma^{0,0}_{0,0}=1$ it follows that $x^{0,0}_{0,0}=1$.
(ii) Remark 12 shows that, given $E_{x},E^{\dagger}_{y}\in\mathcal{E}_{n}$, the value $t-p$ counts the number of places in which a nontrivial Pauli element in $E_{x}$ does not cancel a nontrivial Pauli element in $E^{\dagger}_{y}$. As a consequence, if $t-p$ is even, then $E_{x}$ and $E^{\dagger}_{y}$ commute; if $t-p$ is odd, then $E_{x}$ and $E^{\dagger}_{y}$ anticommute. Since $\Gamma_{xy}=0$ if $E_{x}E^{\dagger}_{y}+E^{\dagger}_{y}E_{x}=0$, Eq. (135) then leads to $x^{t,p}_{i,j}=0$ if $t-p$ is odd. This proves condition (ii).
@@
$\Gamma_{xy}=\Gamma_{x^{\prime}y^{\prime}}\,,\quad\quad\quad$ $\text{if}\quad\quad E_{x}E^{\dagger}_{y}-E^{\dagger}_{y}E_{x}=0\quad\text{and}$
@@
Recall from Remark 12 that $\operatorname{wt}(E_{x}E^{\dagger}_{y})=i+j-t-p$ and note that $t-p=t^{\prime}-p^{\prime}$ for any permutation from Eq. (8.3). Therefore if $t-p=t^{\prime}-p^{\prime}$ is even and $(i,j,i+j-t-p)$ is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$, then
@@
Eq. (135) shows that $\gamma^{0,0}_{j,0}x^{0,0}_{j,0}=\sum_{x\in{\boldsymbol{4}}^{n},\,\operatorname{wt}(E_{x})=j}\Gamma_{x0}$. This allows us to write Eq. (63) as Eq. (140). By Eq. (135), the condition for pure codes stated in Eq. (66) can be written as $x^{t,p}_{i,j}=0$ for $0<i<\delta$ or $0<j<\delta$ or $0<i+j-t-p<\delta$. But this is nothing else than Eq. (141). This ends the proof. ∎
@@
Eq. (64) shows that stabilizer codes satify $0\leq\Gamma_{xy}\leq\Gamma_{xx}$ Recall that $\tilde{\Gamma}\propto\sum_{\pi\in\operatorname{Aut}(4,n)}\pi(\Gamma)$ [see Eq. (109)]. Thus the same relation holds after averaging and $0\leq\tilde{\Gamma}_{xy}\leq\tilde{\Gamma}_{xx}$ since $\pi$ permutes rows and columns of $\Gamma$.
Because $\tilde{\Gamma}$ is spanned by $M_{i,j}^{t,p}$, which have entries in $\{0,1\}$ is follows that $0\leq x^{t,p}_{i,j}\leq{x^{i,i}_{i,i}}$. According to condition (iii) from Proposition 18, $x^{i,i}_{i,i}=x^{0,0}_{i,0}$ and therefore, $0\leq x^{t,p}_{i,j}\leq x^{0,0}_{i,0}$. ∎
@@ 8.4. Symmetry-reduced SDP bound
We now state the symmetry reduced version of the main semidefinite programming bound in Eq. (5.7). It includes all constraints on $x^{t,p}_{i,j}$ derived in the previous sections [Propositions 16 to 19].
@@
    The shadow inequalities [see Eq. (18)] leads to
@@
    with equality for $K=1$ and $n-j$ odd [57, Theorem 12].
@@
    (145) $\sum_{j\geq 0}\gamma_{2j,0}^{0,0}x_{2j,0}^{0,0}=\begin{cases}2^{n-\log_{2}(K)-1}\quad\text{(Type I)\,,}\\ 2^{n-\log_{2}(K)}\quad\quad\!\text{(Type II)\,.}\end{cases}$
@@
We conclude this section by point out that the main semidefinite bound in Eq. (5.7) with scaling of $O(4^{n})$ is now symmetry reduced to Eq. (21) with scaling of $O(n^{4})$. This fact allows us to handle codes with larger $n$.
@@ 9.1. The Lovász bound refutes a $(\!(7,1,4)\!)_{2}$ code
Using the feasibility version of the symmetry reduced Lovász number SDP for self-dual quantum codes [see Eq. (5.8)], we prove the non-existence of a $(\!(7,1,4)\!)_{2}$ code. The feasibility version includes as an extra condition the constraint $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{xx}=2^{n}$:
@@
Following Section 8.4 we symmetry reduce the SDP (9.1) as done for the main semidefinite programming bound [Eq. (5.7)]. First apply conditions (i) and (iii) from Proposition (18) to map $\Gamma_{00}=1$ and $\Gamma_{x0}=\Gamma_{xx}$ to $x^{0,0}_{i,0}$ and $x^{0,0}_{i,0}=x^{i,i}_{i,i}$, respectively. Then consider the block-diagonalization of $\tilde{\Gamma}\succeq 0$ as done with $\tilde{R}$ [Eq. (85)]. Recall that $\Gamma_{xy}=0$ if $x\sim y$. This implies that $x^{t,p}_{i,j}=0$ if $t-p$ is odd or $\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset$. Here we have used the condition for pure codes [Eq. (141)] and anti-commutator condition (ii) of Proposition (18). Finally, $\operatorname{tr}(\Gamma)=1$ is equivalent to $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{x0}=2^{n}$ due to $\Gamma_{x0}=\Gamma_{xx}$. This implies $\sum^{n}_{i=0}x^{0,0}_{i,0}\gamma^{0,0}_{i,0}=\frac{2^{n}}{K}$ by condition (iv) in Proposition 18. Thus the symmetry reduced version of Eq. (9.1) reads:
@@
The dual problem can be used to prove infeasibility of a primal problem [see Eq. (37)]. Appendix A, Proposition 25 shows that the dual program is:
@@
$y^{t,p}_{i,j}=0\hskip 80.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
@@
where the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$.
If $\alpha>0$, then there do not exist a set of $x^{t,p}_{i,j}$ satisfying the constraints from Eq. (9.1). This is the case for the $(\!(7,1,4)\!)_{2}$ code:
Observation 23.
@@
9.2. The symmetry-reduced SDP refutes $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes
We first remark the following: in principle, adding constraints in the primal program makes a semidefinite program stronger. From a numerical perspective however this can be disadvantageous: The more constraints appear in the primal program, the larger is the span of the dual feasible region. As a consequence, the dual program can numerically be more costly to solve. For obtaining a nice infeasibility certificate it can thus help to drop a number of constraints.
For refuting the $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ code, the following relaxation of the SDP (21) of Theorem 21 is enough:
@@ 9.2. The symmetry-reduced SDP refutes $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). The SDP (9.2) is obtained from SDP 21 by removing the constraint corresponding to $\tilde{\Gamma}^{c}\succeq 0$ and to only take the equality conditions in $K2^{-n}\sum^{n}_{i=0}K_{j}(i,n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\,\geq\,\gamma^{0,0}_{j,0}x^{0,0}_{j,0}$.
@@
Here the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
Observation 24.
The SDP from Eq. (9.2) has solutions with $\alpha>0$ for the parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$. In particular, for the $(\!(8,9,3)\!)_{2}$, there is a dual solution with $\alpha\approx 0.989$, maximum constraint violation of $\mathcal{O}(10^{-13})$, and minimum eigenvalue $\lambda_{\min}\approx 0.0010$. For $(\!(10,5,4)\!)_{2}$, there is a dual solution with $\alpha\approx 1$, maximum constraint violation of $\mathcal{O}(10^{-13})$ and minimum eigenvalue $\lambda_{\min}\approx 0.050$. As a consequence, codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ do not exist.
9.3. Code availability
To solve the SDPs from Eq. (9.1) and Eq. (9.2), we used the Python API PICOS [62] toghether with the SDP solver CVXOPT [43]. The program can be found online in https://github.com/ganglesmunne/SDP_bounds_for_quantum_codes. To obtain solutions with some rational values we run the SDP iteratively, where in each iteration an an entry of $Y^{(a,k)}$ was forced to be equal to a rounded previous entry. This procedure leads to the rational solution in Appendix B.
@@ 10. Discussion
We provided a complete semidefinite programming hierarchy for the existence of quantum codes, and relaxations that recover a Lovász bound for self-dual quantum codes, a quantum Delsarte bound, and the quantum version of Schrijver and Gijswijt’s SDP bounds that use the Terwilliger algebra. With this we show that $(\!(7,1,4)\!)_{2}$, $(\!(8,9,3)\!)_{2}$, and $(\!(10,5,4)\!)_{2}$ codes do not exist. While the non-existence of $(\!(7,1,4)\!)_{2}$ was known previously through analytical methods [30], to our knowledge the latter two bounds seem to be new [61, Table III].
An interesting aspect are the changes that appear in the quantum versions of these semidefinite programming bounds. These are perhaps best illustrated in the Lovász bound [Eqs. (3.5) and (3.5)]: in the classical case the Hamming distance indicates which entries in $M$ are set to zero. For quantum codes also the anti-commutation relations between pairs of observables yield constraints: $\Gamma_{ab}$ is additionally set to zero if the Pauli strings $E_{a}$, $E_{b}$ anticommute. This can be captured by a graph with loops, and makes the quantum Lovász bound naturally fit into a recently developed framework for a hypergraph Lovász $\vartheta$ number [10]. In particular, this leads to both different quantum Lovász and quantum Delsarte bounds compared to classical ${\bf{4}}^{n}$ codes.
@@
We note that our hierarchy also applies to the existence of absolutely maximally entangled and $m$-uniform states, these corresponding to self-dual (pure) codes [60, 28]. With small modifications, the hierarchy also lower bounds the average purities of $n$-qubit states, such as found in the Meyer-Wallach [45, 8, 66], potential of multipartite entanglement [19], and concentratable entangled measures [63].
@@
  • The shadow bounds provide powerful constraints, e.g. these alone refute the existence of a $(\!(4,1,3)\!)_{2}$ code. Is there a formulation of the generalized shadow bounds that can be useful as part of an SDP hierarchy [58]?
@@
Finally, our approach is formally dimension-free. That is, on can target scenarios where the Hilbert space is not known. This could be interesting for the study of quantum version of the quantum zero-error communication capacity of a quantum graph [17].

### v2

@@ 8.2. Stabilizer codes
Consider the case of a stabilizer code with stabilizer group $S$. Then we can recover the Propositions 17 and 18 in an alternative fashion for the matrices
@@
that is, by replacing $Q=\{0^{n}\}$ with $Q=\mathcal{S}$ where $\mathcal{S}$ contains the unsigned Pauli strings of $S$ mapped to $4^{n}$.
Then as in the classical case discussed in Sections 6.3 and 6.4, these matrices are both positive semidefinite and invariant under $\operatorname{Aut}_{0}(4,n)$. They can thus be expressed in a basis of the Terwilliger algebra $\mathcal{A}_{4,n}$, recovering the same constraints on $x^{t,p}_{i,j}$ as in the general case. This follows from the fact that by the definition of $\Gamma^{v}$ in Eq. (96), it holds that $\Gamma^{v}=\Gamma^{0}$ for any $v\in\mathcal{S}$ leading to $\tilde{\Gamma}=\Gamma^{0}$, and that for a stabilizer code $|S|=|\mathcal{S}|=2^{n}/K$.
@@
which recovers the general case [Eq. (111)]. In analogy to Proposition (18),
@@
(132) $\tilde{\Gamma}^{c}(\mathcal{S})$ $=\frac{|\mathcal{S}|}{4^{n}-|\mathcal{S}|}\sum_{(i,j,t,p)\in\mathcal{I}(4,n)}(x^{0,0}_{i+j-t-p}-x^{t,p}_{i,j})M^{t,p}_{i,j}\,$,
@@ 8.3. Symmetry-reduced constraints
Recall from Eq. (123) that
@@
(i) From $\Gamma_{00}=1$, $(M_{0,0}^{0,0})_{00}=1$ and is zero otherwise, and from $\gamma^{0,0}_{0,0}=1$ it follows that $x^{0,0}_{0,0}=1$.
(ii) Remark 13 shows that, given $E_{x},E^{\dagger}_{y}\in\mathcal{E}_{n}$, the value tpt-p counts the number of places in which a nontrivial Pauli element in $E_{x}$ does not cancel a nontrivial Pauli element in $E^{\dagger}_{y}$. As a consequence, if tpt-p is even, then $E_{x}$ and $E^{\dagger}_{y}$ commute; if tpt-p is odd, then $E_{x}$ and $E^{\dagger}_{y}$ anticommute. Since $\Gamma_{xy}=0$ if $E_{x}E^{\dagger}_{y}+E^{\dagger}_{y}E_{x}=0$, Eq. (134) then leads to $x^{t,p}_{i,j}=0$ if tpt-p is odd. This proves condition (ii).
@@
$\Gamma_{xy}=\Gamma_{x^{\prime}y^{\prime}}\,$$\quad\quad\quad$ $\text{if}\quad\quad E_{x}E^{\dagger}_{y}-E^{\dagger}_{y}E_{x}=0\quad\text{and}$
@@
Recall from Remark 13 that $\operatorname{wt}(E_{x}E^{\dagger}_{y})=i+j-t-p$ and note that $t-p=t^{\prime}-p^{\prime}$ for any permutation from Eq. (8.3). Therefore if $t-p=t^{\prime}-p^{\prime}$ is even and (i,j,i+jtp)(i,j,i+j-t-p) is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$, then
@@
Eq. (134) shows that $\gamma^{0,0}_{j,0}x^{0,0}_{j,0}=\sum_{x\in{\boldsymbol{4}}^{n},\,\operatorname{wt}(E_{x})=j}\Gamma_{x0}$. This allows to write Eq. (66) as Eq. (139). By Eq. (134), the condition for pure codes stated in Eq. (68) can be written as $x^{t,p}_{i,j}=0$ for $0<i<\delta$ or $0<j<\delta$ or $0<i+j-t-p<\delta$. But this is nothing else than Eq. (140). This ends the proof. ∎
@@
Eq. (67) shows that stabilizer codes satify $0\leq\Gamma_{xy}\leq\Gamma_{xx}$. Recall that $\tilde{\Gamma}\propto\sum_{\pi\in\operatorname{Aut}(4,n)}\pi(\Gamma)$ [see Eq. (111)]. Thus the same relation holds after averaging and $0\leq\tilde{\Gamma}_{xy}\leq\tilde{\Gamma}_{xx}$ since $\pi$ permutes rows and columns of $\Gamma$.
Because $\tilde{\Gamma}$ is spanned by $M_{i,j}^{t,p}$, which have entries in $\{0,1\}$, is follows that $0\leq x^{t,p}_{i,j}\leq{x^{i,i}_{i,i}}$. According to condition (iii) from Proposition 19, $x^{i,i}_{i,i}=x^{0,0}_{i,0}$ and thus $0\leq x^{t,p}_{i,j}\leq x^{0,0}_{i,0}$. ∎
@@ 8.4. Symmetry-reduced SDP bound
We now state the symmetry reduced version of the main semidefinite programming bound in Eq. (5.7). It includes all constraints on $x^{t,p}_{i,j}$ derived in the previous sections [Propositions 17 to 20].
@@
    The shadow inequalities [Eq. (18)] leads to
@@
    with equality for $K=1$ and njn-j odd [62, Theorem 12].
@@
[…]
@@
We conclude this section by pointing out that the main semidefinite bound in Eq. (5.7) with scaling of $O(4^{n})$ is now symmetry reduced to Eq. (22) with scaling of $O(n^{4})$. This fact allows us to handle codes with larger $n$.
@@ 9.1. The Lovász bound refutes a $(\!(7,1,4)\!)_{2}$ code
Using the feasibility version of the symmetry reduced Lovász theta number SDP for self-dual quantum codes [Eq. (5.8)], we prove the non-existence of a $(\!(7,1,4)\!)_{2}$ code. The feasibility version includes as an extra condition the constraint $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{xx}=2^{n}$:
@@
Following Section 8.4 we symmetry reduce the SDP in Eq. (9.1) as done for the main semidefinite programming bound [Eq. (5.7)]. First apply conditions (i) and (iii) from Proposition (19) to map $\Gamma_{00}=1$ and $\Gamma_{x0}=\Gamma_{xx}$ to $x^{0,0}_{i,0}$ and $x^{0,0}_{i,0}=x^{i,i}_{i,i}$, respectively. Then consider the block-diagonalization of $\tilde{\Gamma}\succeq 0$ as done with $\tilde{R}$ [Eq. (87)]. Recall that $\Gamma_{xy}=0$ if $x\sim y$. This implies that $x^{t,p}_{i,j}=0$ if tpt-p is odd or $\{i,j,i+j-t-p\}\cap\{1,\cdots,\delta-1\}\neq\emptyset$. Here we have used the condition for pure codes [Eq. (140)] and anti-commutator condition (ii) of Proposition (19). Finally, $\operatorname{tr}(\Gamma)=1$ is equivalent to $\sum_{x\in{\boldsymbol{4}}^{n}}\Gamma_{x0}=2^{n}$ due to $\Gamma_{x0}=\Gamma_{xx}$. This implies $\sum^{n}_{i=0}x^{0,0}_{i,0}\gamma^{0,0}_{i,0}=\frac{2^{n}}{K}$ by condition (iv) in Proposition 19. Thus the symmetry reduced version of Eq. (9.1) reads:
@@
Weak duality allows to prove infeasibility of a primal problem from its dual. Appendix A, Proposition 26 shows that the program dual to that of Eq. (27) is:
@@
[…]
@@
Here the dual variables are real matrices $Y^{(a,k)}$ of size n+a2kn+a-2k with $0\leq a\leq k\leq n+a-k$.
If $\alpha>0$, then there do not exist a set of $x^{t,p}_{i,j}$ satisfying the constraints from Eq. (9.1), and thus no corresponding code exists. This is the case for the $(\!(7,1,4)\!)_{2}$ code:
Observation 24.
@@
9.2. SDP refutations for the $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes
We first remark the following: in principle, adding constraints in the primal program makes a semidefinite program stronger. From a numerical perspective however this can be disadvantageous: the more constraints appear in the primal program, the larger is the span of the dual feasible region. As a consequence, the dual program can numerically be more costly to solve. For obtaining a nice infeasibility certificate it can thus help to reduce the number of constraints.
For refuting the $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes, the following relaxation of the SDP in Eq. (22) of Theorem 22 suffices:
@@ 9.2. SDP refutations for the $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). The SDP of Eq. (9.2) is obtained from the SDP in Eq. 22 by removing the constraint corresponding to $\tilde{\Gamma}^{c}\succeq 0$ and to only take the equality conditions in $K2^{-n}\sum^{n}_{i=0}K_{j}(i,n)\gamma^{0,0}_{i,0}x^{0,0}_{i,0}\,\geq\,\gamma^{0,0}_{j,0}x^{0,0}_{j,0}$.
@@
Here the dual variables are real matrices $Y^{(a,k)}$ of size n+a2kn+a-2k with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
Observation 25.
The SDP from Eq. (9.2) has solutions with $\alpha>0$ for the parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$. In particular, for the $(\!(8,9,3)\!)_{2}$, there is a dual solution with $\alpha\approx 0.989$, maximum constraint violation of $\mathcal{O}(10^{-13})$, and minimum eigenvalue $\lambda_{\min}\approx 0.0010$. For $(\!(10,5,4)\!)_{2}$, there is a dual solution with $\alpha\approx 1$, maximum constraint violation of $\mathcal{O}(10^{-13})$ and minimum eigenvalue $\lambda_{\min}\approx 0.050$. As a consequence, codes with parameters $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ do not exist. The infeasibility certificates can be found online at https://github.com/ganglesmunne/SDP_bounds_for_quantum_codes [Ref. [48]].
@@ 10. Discussion
We provided a complete semidefinite programming hierarchy for the existence of quantum codes, and relaxations that recover a Lovász bound for self-dual quantum codes, a quantum Delsarte bound, and the quantum version of Schrijver and Gijswijt’s semidefinite programming bounds that use the Terwilliger algebra. With this we show that $(\!(7,1,4)\!)_{2}$, $(\!(8,9,3)\!)_{2}$, and $(\!(10,5,4)\!)_{2}$ codes do not exist. While the non-existence of $(\!(7,1,4)\!)_{2}$ was known previously through analytical methods [31], to our knowledge the latter two bounds are new [65, Table III].
An interesting aspect are the differences between the classical and quantum versions of these semidefinite programming bounds. These are perhaps best illustrated in the Lovász bound [Eqs. (3.5),  (3.5)]: in the classical case the Hamming distance indicates which entries in $M=\chi\chi^{T}$ are set to zero. For quantum codes the anti-commutation relations between pairs of observables yield additional constraints: $\Gamma_{ab}$ is set to zero if the Pauli strings $E_{a}$, $E_{b}$ anticommute, which can be captured by a graph with loops. In particular, this leads to different quantum Lovász and quantum Delsarte bounds than in the classical case of quaternary codes.
@@
We note that our hierarchy also applies to the existence of absolutely maximally entangled and $m$-uniform states, these corresponding to self-dual codes [64, 29]. With small modifications, the hierarchy also lower bounds the average purities of $n$-qubit states, such as found in the Meyer-Wallach [46, 8, 70], potential of multipartite entanglement [19], and concentratable entangled measures [67].
Finally, we want to highlight that our approach is formally dimension-free. That is, on can target scenarios where the Hilbert space is not known. This could be interesting for the study of quantum version of the quantum zero-error communication capacity of a quantum graph [17].
@@
  • The shadow bounds provide powerful constraints, e.g., these alone refute the existence of a $(\!(4,1,3)\!)_{2}$ code. Is there a formulation of the generalized shadow bounds that can be useful as part of an SDP hierarchy [63]?
  • Does our main SDP relaxation imply the shadow bounds? If not, are they included at some finite level of the hieararchy?
@@
10.1. Related work
Lai et al. [39] have introduced an SDP relaxation in the context of entanglement-assisted codeword stabilized codes. When restricted to the case with no pre-shared entanglement, their bounds showed no advantage over the LP bounds. Berta et al. [6] presented a SDP hierarchy to determine the existence of a quantum code based on the symmetric extension hierarchy [15]; see also the follow up work [38]. This hierarchy is also complete but does not easily allow for numerically tractable relaxations. In the classical case, a complete hierarchy for the existence of codes can be formulated using 010-1 programming for maximum independent sets [44, 40].
Data and code availability
To solve the SDPs from Eq. (9.1) and Eq. (9.2), we used the Python API PICOS [66] together with the SDP solver CVXOPT [10]. The program and the SDP solutions can be found online in https://github.com/ganglesmunne/SDP_bounds_for_quantum_codes [Ref. [48]]. To obtain solutions with some rational values we run the SDP iteratively, where in each iteration an an entry of $Y^{(a,k)}$ was forced to be equal to a rounded previous entry. This procedure leads to the rational solution in Appendix B.
Conflict of interest
The authors declare no conflict of interest.